# Jason Zhaoxin Fan3, Bo Tang2, Shichao Song1, Mengwei Wang1, Jiawei Yang1

### v1

@@ 5.4.2 Selection of Evaluator
Inter-context Conflict 99.10 95.65
@@
As shown in Table 2, We present the evaluation metrics and their consistency with human judgments. The ASR and AFR metric exhibit a high human consistency. Similarly, the F1 (correct) and F1 (incorrect) scores obtained using DeepSeek also demonstrate strong agreement with human judgments. Therefore, DeepSeek is uniformly adopted for evaluation across all experiments.
@@ 6 Conclusion
This paper introduces SafeRAG, a benchmark designed to assess the security vulnerabilities of RAG against data injection attacks. We identified four critical attack surfaces: noise, conflict, toxicity, and DoS, and revealed significant weaknesses across the retriever, filter, and generator components of RAG. By proposing novel attack strategies such as silver noise, inter-context conflict, soft ad, and white DoS, we exposed critical gaps in existing defenses and demonstrated the susceptibility of RAG systems to subtle yet impactful threats.
@@ Appendix A Appendix
A.1 Preliminaries and Definitions
A.1.1 RAG Pipeline
@@
A.1.2 SafeRAG Dataset Construction
(1) Base Dataset Construction: Raw document is collected from the external news website, and paragraphs meeting the following criteria are selected: 1) contain more than 8 consecutive sentences; 2) consecutive sentences revolve around a specific topic; 3) consecutive sentences can generate comprehensive questions of the what, why, or how types. For each paragraph, a comprehensive question $q_{j}$ is generated, and $n$ golden contexts $C_{g}^{j}=\{c_{g}^{1},c_{g}^{2},\dots,c_{g}^{n}\},n\geq 8$ closely related to the question $q_{j}$ are extracted. Each golden context $c_{g}^{j}\in C_{g}^{j}$ is manually screened and verified to ensure accuracy, coherence, and the ability to fully answer the question. The base dataset is defined as:
@@ A.1.2 SafeRAG Dataset Construction
where $i$ represents the uniform instruction $i$ provided to the LLM [^8] The default instruction $i$ used in this paper is shown in Fig. 14.
(2) Attacking Texts Generation and RAG Security Evaluation Dataset Construction: To evaluate the security of RAG under different adversarial scenarios, we design four attack tasks $T=\{\mathrm{SN},\mathrm{ICC},\mathrm{SA},\mathrm{WDoS}\}$: silver noise $\mathrm{SN}$, inter-context conflict $\mathrm{ICC}$, soft ad $\mathrm{SA}$, and white DoS $\mathrm{WDoS}$. For each attack task $t\in T$, malicious attack texts are generated and combined with golden contexts to construct the RAG security evaluation dataset (i.e., SafeRAG dataset). The detailed process is as follows:
@@
The complete SafeRAG dataset is defined as:
@@
A.1.3 Threat Framework: Attacks on the RAG Pipeline
@@
Regardless of the stage at which the injection occurs, the attacker’s objective is to mislead or compromise the RAG pipeline’s final output $r$ by leveraging the attack contexts $C_{a}^{t,j}$. It is important to note that, under the attack assumptions in this paper, golden contexts $C_{g}^{j}$ are neither altered in content nor re-ranked [^10] This setting is widely adopted in numerous many attacks Xiang et al. (2024); Zhong et al. (2023); Zou et al. (2024); Pan et al. (2023b, a); Du et al. (2022). Through this method, the attacker maximizes the system’s original usability and normalcy while covertly influencing the RAG pipeline’s generated responses.
@@
Model Noise Conflict Toxicity DoS
F1_Variants AFR F1_Variants AFR F1_Variants AFR F1_Variants AFR
DeepSeek 0.1032 - 0.4000 0.69 - 0.38 0.2068 0.17
GPT-3.5-Turbo 0.1102 - 0.3615 0.72 - 0.47 0.1654 0.17
GPT-4 0.1141 - 0.3615 0.66 - 0.29 0.4760 0.16
GPT-4o 0.1229 - 0.3615 0.68 - 0.37 0.3196 0.28
Qwen 7B 0.1016 - 0.4948 0.75 - 0.47 0.2582 0.57
Qwen 14B 0.1005 - 0.5000 0.65 - 0.38 0.1842 0.29
Baichuan 13B 0.0706 - 0.4800 0.87 - 0.59 0.7222 1.00
ChatGLM 6B 0.0815 - 0.5966 0.55 - 0.27 0.5096 1.00

### v2

@@ 5.4.2 Selection of Evaluator
Inter-context Conflict 99.10 98.48 95.65
@@
As shown in Table 2, We present the evaluation metrics and their consistency with human judgments. The ASR/AFR metric exhibits a high human consistency. Similarly, the F1 (correct) and F1 (incorrect) scores obtained using DeepSeek also demonstrate strong agreement with human judgments. Therefore, DeepSeek is uniformly adopted for evaluation across all experiments.
@@ 6 Conclusion
This paper introduces SafeRAG, a benchmark designed to assess the security vulnerabilities of RAG against data injection attacks. We identified four critical attack tasks: noise, conflict, toxicity, and DoS, and revealed significant weaknesses across the retriever, filter, and generator components of RAG. By proposing novel attack strategies such as silver noise, inter-context conflict, soft ad, and white DoS, we exposed critical gaps in existing defenses and demonstrated the susceptibility of RAG systems to subtle yet impactful threats.
@@ Appendix A Appendix
A.1 Limitations
Despite the comprehensive evaluation framework provided by SafeRAG, there are still some limitations to be addressed:
(1) Attack Coverage: SafeRAG primarily focuses on data injection attacks, assessing vulnerabilities in RAG pipeline. It does not evaluate other orthogonal security threats, such as model backdoor attacks, which could compromise RAG at the model level. Future work could extend SafeRAG to incorporate a broader range of security risks beyond data manipulation.
(2) Modal Limitations: SafeRAG primarily targets single-modal, unstructured textual RAG, without considering multimodal RAGs that integrate images, structured graphs, or audio for retrieval and generation. Given the growing adoption of multimodal RAG, future work should explore SafeRAG’s adaptation to multimodal and structured knowledge retrieval scenarios.
Despite these limitations, SafeRAG provides the first Chinese security benchmark for RAG, offering valuable insights into their robustness against data injection attacks and laying the groundwork for future security research.
A.2 Potential Risks
By standardizing a security evaluation framework, SafeRAG may inadvertently assist adversaries in understanding how RAG is tested, allowing them to designmore evasive attack strategies that exploit weaknesses beyond the scope of current evaluations. This highlights the need for continuously updating attack methodologies and expanding RAG security evaluation techniques.
A.3 Ethical Considerations
SafeRAG does not collect or use personally identifiable information (PII) or offensive content. The dataset is built from publicly available news articles and open-source knowledge bases, explicitly excluding any sensitive or restricted sources. Designed for academic and security research, SafeRAG focuses on improving RAG robustness without involving user-generated or proprietary data. By adhering to strict data integrity and ethical standards, SafeRAG ensures a responsible and secure benchmark for evaluating RAG security.
A.4 Preliminaries and Definitions
A.4.1 RAG Pipeline
@@
A.4.2 Base Dataset Construction
Raw document is collected from the external news website, and paragraphs meeting the following criteria are selected: 1) contain more than 8 consecutive sentences; 2) consecutive sentences revolve around a specific topic; 3) consecutive sentences can generate comprehensive questions of the what, why, or how types. For each paragraph, a comprehensive question $q_{j}$ is generated, and $n$ golden contexts $C_{g}^{j}=\{c_{g}^{1},c_{g}^{2},\dots,c_{g}^{n}\},n\geq 8$ closely related to the question $q_{j}$ are extracted. Each golden context $c_{g}^{j}\in C_{g}^{j}$ is manually screened and verified to ensure accuracy, coherence, and the ability to fully answer the question. The base dataset is defined as:
@@ A.4.2 Base Dataset Construction
where $i$ represents the uniform instruction $i$ provided to the LLM [^19] The default instruction $i$ used in this paper is shown in Fig. 14.. The distribution of the base dataset is shown in Table 3.
Table 3: Distribution of the base dataset.
Domain What How Why
Politics 10 7 2
Finance 10 7 3
Technology 12 5 2
Culture 7 13 1
Military 12 3 6
The politics and finance domains exhibit a similar pattern, with a higher number of what and how questions compared to why questions. This is because news coverage in these areas primarily focuses on reporting events, policies, and market trends, which naturally correspond to what questions (e.g., What policies were introduced? What were the market movements?). How questions are also relatively frequent, as they are used to explain processes and mechanisms (e.g., How does a new financial regulation impact the market?). In contrast, why questions are less common, as political and financial reporting tends to present facts rather than analyze motivations, leaving deeper interpretation to opinion pieces or expert analyses.
In the technology domain, what questions dominate, given that news in this field often revolves around new products, scientific advancements, and industry developments (e.g., What is the latest AI breakthrough?). While some how questions appear in discussions of technological mechanisms and implementations, why questions are rare, as most technology reporting focuses on descriptive rather than explanatory narratives.
The culture domain exhibits a distinct pattern, with how questions being the most frequent. Cultural discussions often revolve around trends, artistic movements, and societal changes, which naturally lead to explanations of processes (e.g., How has digital art influenced modern design?). In contrast, what questions are fewer, as cultural reporting tends to be less event-driven, and why questions are extremely rare, given that cultural phenomena are often subjective and interpretative rather than objective and causal.
For military topics, the data shows a relatively high proportion of why questions, second only to what questions. This can be attributed to the nature of military reporting, which often involves analyzing strategic decisions, conflicts, and security developments (e.g., Why did a country conduct military drills?). what questions are still the most common, given that military news frequently reports events, operations, and technological advancements, while how questions appear less frequently, as military strategies and tactics are often classified.
A.4.3 SafeRAG Dataset Construction
To evaluate the security of RAG under different adversarial scenarios, we design four attack tasks $T=\{\mathrm{SN},\mathrm{ICC},\mathrm{SA},\mathrm{WDoS}\}$: silver noise $\mathrm{SN}$, inter-context conflict $\mathrm{ICC}$, soft ad $\mathrm{SA}$, and white DoS $\mathrm{WDoS}$. For each attack task $t\in T$, malicious attack texts are generated and combined with golden contexts to construct the RAG security evaluation dataset (i.e., SafeRAG dataset). The detailed process is as follows:
@@ A.4.3 SafeRAG Dataset Construction
The complete SafeRAG dataset is defined as [^20] All manual annotation tasks in the dataset construction process were conducted by professionals with a background in journalism, ensuring high-quality annotations. Additionally, the gender ratio of the annotators was balanced at 1:1.:
@@
Model Noise Conflict Toxicity DoS
F1_Variants AFR F1_Variants AFR F1_Variants AFR F1_Variants AFR
DeepSeek 0.1032 - 0.4000 0.69 - 0.38 0.2068 0.17
GPT-3.5-Turbo 0.1102 - 0.3615 0.72 - 0.47 0.1654 0.17
GPT-4 0.1141 - 0.3615 0.66 - 0.29 0.4760 0.16
GPT-4o 0.1229 - 0.3615 0.68 - 0.37 0.3196 0.28
Qwen 7B 0.1016 - 0.4948 0.75 - 0.47 0.2582 0.57
Qwen 14B 0.1005 - 0.5000 0.65 - 0.38 0.1842 0.29
Baichuan 13B 0.0706 - 0.4800 0.87 - 0.59 0.7222 1.00
ChatGLM 6B 0.0815 - 0.5966 0.55 - 0.27 0.5096 1.00
A.4.4 Threat Framework: Attacks on the RAG Pipeline
@@
Regardless of the stage at which the injection occurs, the attacker’s objective is to mislead or compromise the RAG pipeline’s final output $r$ by leveraging the attack contexts $C_{a}^{t,j}$. It is important to note that, under the attack assumptions in this paper, golden contexts $C_{g}^{j}$ are neither altered in content nor re-ranked [^22] This setting is widely adopted in numerous many attacks Xiang et al. (2024); Zhong et al. (2023); Zou et al. (2024); Pan et al. (2023b, a); Du et al. (2022).. Through this method, the attacker maximizes the system’s original usability and normalcy while covertly influencing the RAG pipeline’s generated responses.
@@
Figure 28: Filter NLI.
Figure 29: Compressor SKR.
Figure 30: The data format of SafeRAG dataset.